# Towards Training-Free Open-World Classification with 3D Generative Models

### v1

@@ 3.4 Inference Phase
$\hat{TS_{m}}=\frac{\mathbf{FPS}(TS_{m},\hat{N_{p}})-\mathbf{Avg}(TS_{m})}{\mathbf{L}(TS_{m})}\cdot R,\quad R\in\mathbb{R}^{3\times 3},$ (5)
@@
$TF_{m}=\mathbf{RMs}\left(\hat{TS_{m}}\right),\quad TF_{m}\in\mathbb{R}^{\hat{N_{p}}\times D}.$ (6)
@@
4 Experiment
4.1 Experiment Setup
Datasets. The effectiveness of the proposed method is validated using the public datasets ModelNet40, ModelNet10, and McGill, which were introduced by OP3D Zhao et al. (2025).
@@ 4.1 Experiment Setup
ModelNet40/10 evolves from ModelNet40/10 Wu et al. (2015) through random rotation of the samples, resulting in open-pose variations. ModelNet40 is a benchmark dataset in the field of 3D point cloud classification. It is derived from computer-aided design models and contains high-quality 3D synthetic objects, free from noise and with regular structures. The dataset covers 40 categories of everyday objects (such as pianos, tables, airplanes, etc.) and consists of 12,311 3D objects, each represented by 10,000 points. The points are represented by 3D coordinates (x, y, z) and may include additional information such as normal vectors or colors (R, G, B). ModelNet10 is a curated subset of ModelNet40, focusing on 10 common object categories, such as beds, monitors, and chairs, containing a total of 4,899 3D objects.
@@
McGill Zhao et al. (2025) is derived from McGill Siddiqi et al. (2008) by randomly rotating the samples in McGill and removing five categories that overlap with ModelNet40. McGill includes 14 categories of complex geometric shapes, such as ants, snakes, and hands. The objects in this dataset exhibit higher morphological diversity, particularly posing greater challenges when handling natural object categories. Each sample consists of point coordinates (x, y, z) and topological information.
Evaluation Metric. Following SOTAsZhang et al. (2023); Huang et al. (2023); Zhao et al. (2025), we evaluate the open-world classification performance using top-1 accuracy and mean accuracy (mAcc). Top-1 accuracy measures the proportion of samples for which the model’s top prediction matches the ground truth, while mAcc calculates the average accuracy across all categories to assess the balance of performance among classes.
Implement Details. All feature extraction and inference tasks were performed on an NVIDIA RTX 3090 GPU. For the point cloud diffusion model, we utilize a random seed $s\sim\mathcal{U}(0,50)$ to introduce diversity into the generated anchor point clouds of any category.
During the data augmentation stage, all the anchor samples and test samples are downsampled to 1024 points by the farthest point sampling. For the feature output dimensions of different models, the TET pre-trained model outputs features with dimensions of (256, 1), and the TAP pre-trained model outputs features with dimensions of (256, 384). To facilitate subsequent feature comparison and processing, we apply max pooling along the second dimension of the TAP model’s output, ultimately unifying the features to dimensions of (256, 1).
4.2 Performance Comparison
Our approach is compared against six SOTA methods: PointCLIP Zhang et al. (2023), ULIP Xue et al. (2023), ReconCLIP Qi et al. (2023), CLIP2Point Huang et al. (2023), PointCLIPv2 Zhu et al. (2023), and OP3D Zhao et al. (2025), on the open-pose and open-category classification datasets ModelNet10 and McGill. To ensure fairness, we utilized the pre-trained models from the official GitHub repositories provided by the respective papers (TAP and TET) and strictly followed the original experimental settings. The results presented here do not include the samples generated from LLM-generated prompts. That is, all anchor samples were directly generated using the category names. The impact of LLM-generated prompts on the results will be discussed in subsequent sections.
Our approach exhibits notable enhancements in classification accuracy over current models in both open-pose and open-category recognition tasks. The performance of our model on ModelNet10 dataset is shown in Table 1. Notably, we achieved a 32.0% improvement in overall accuracy (oAcc) and a 30.3% improvement in mean accuracy (mAcc). On McGill dataset, our model also demonstrated competitive performance. Compared to the current leading method, our approach improved overall accuracy (oAcc) by 8.7% and mean accuracy (mAcc) by 2.8%.
@@ 4.2 Performance Comparison
Table 1: Comparison to SOTAs. Proj. indicates whether it is necessary to project the 3D samples into 2D images. denotes that there is an overlap between the test categories and the training categories of the pre-trained model utilized.
@@
Ours N 54.6 52.0 47.8 47.5
Figure 3: Comparison on the Number of Anchor Samples $N_{a}$.
Table 2: Comparison of classification effects of anchors generated by different generative models on ModelNet10.
@@
Shap-e 22.0 19.0 66.0 31.4 10.5 65.0 67.4 79.0 65.0 87.0 53.5 51.2
GD 8.0 30.0 61.0 36.0 4.7 64.0 53.5 77.0 61.0 87.0 51.2 48.2
Ensemble 18.0 22.0 73.0 36.0 2.3 65.0 69.8 80.0 65.0 89.0 54.6 52.0
Table 3: Comparison of classification effects of anchors generated by different generative models on McGill. Abbreviations used: Din. = Dinosaur, Dol. = Dolphin, Oct. = Octopus, Qua. = Quadruple, Spe. = Spectacle, Spi. = Spider.
@@
Table 4: Comparison of classification effects of aligned-pose and open-pose settings on the different representation models using Shap-e.
@@
TET 24.0 16.0 69.0 29.1 33.7 69.0 41.8 79.0 67.0 84.0 53.5 51.3
4.3 Analysis of Results
The performance of the proposed method is influenced by several factors, including the number of anchor points, the selection of generative models, and 3D representation models. In this section, we will perform a quantitative analysis to evaluate the impact of each of these factors, thus demonstrating the effectiveness and robustness of our approach. The subsequent discussion will address each factor in detail.
Impact of Anchor Sample Size $N_{a}$.
We examined the impact of varying the number of anchors generated while fixing the dataset, represent model (TET), and generation model to evaluate its effect on classification performance. Three types of anchor generation approaches were tested: Shap-e, GaussianDreamer (GD), and their mixed (Ensemble).
As shown in Fig. 3, as the number of generated samples increased, the classification performance improved and stabilized. However, for the McGill dataset, performance slightly decreased at 8 anchor samples, probably due to individual sample differences causing fluctuations in the smaller anchor size. While more anchor samples generally improve performance, too many can lead to diminishing returns and increased generation time, necessitating a trad-off between performance and efficiency. Based on these findings, we set the number of anchor points to 7 for all the experiments.
Comparison on Generative Models In this experiment, we kept the dataset, represent model (TET), and the number of generated anchor points constant, while only changing the generative models to test their impact on classification performance. Two generative models were selected: Shap-e and GD. Both models can generate point clouds based on the given prompts, but they differ in methodology. Shap-e directly generates a point cloud, whereas GD first generates a coarse point cloud and then optimizes it using 3D Gaussian Splatting. In terms of generation time, Shape required 7 seconds, while GD took 15 minutes. The results of modelnet10 and mcgill are shown in Table 5.
The results on Modelnet10 are shown in Table 2. Overall, Shap-e outperforms GD in terms of classification accuracy, but it does not consistently lead in every category. Both generative models have their strengths: Shap-e excels in some categories, e.g. bathtub, chair, and monitor, while GD performs better in others, e.g. bed, desk, and table. Consequently, combining samples generated by both models results in improved performance. This shows that the diversity of samples generated by 3D generative models significantly affects the performance of the classification model.
Table 5: Performance comparison of classifiers using anchors generated by two 3D generative models, Shap-e and GaussianDreamer (GD), on ModelNet10 and McGill datasets.
Shap-e GD Accuracy
@@ 4.3 Analysis of Results
ModelNet10 53.5 51.2
51.2 48.2
54.6 52.0
McGill 47.0 47.0
39.1 39.0
47.8 47.5
Table 6: Comparison of prompts and the corresponding classification accuracies of samples generated using those prompts for ’dresser’.
Category Text Description $\mathbf{CD}$ oAcc(Dresser)$\uparrow$
A dresser. 10.5
A dresser with five drawers and circular handles. 4.7
An open design dresser with five exposed drawers. 7.0
A small bedside dresser with two drawers”. 12.8
A vintage dresser with carved details and two doors. 18.6
@@
A decorative dresser with floral carvings and four drawers. 93.0
Table 7: Comparison of classification accuracy on the ModelNet10 using Shap-e with Category Name ($\mathbf{CN}$) and ChatGPT ($\mathbf{CD}$) prompts.
@@
$\mathbf{CN}$ 22.0 19.0 66.0 31.4 10.5 65.0 67.4 79.0 65.0 87.0 53.5 51.2
$\mathbf{CD}$ 22.0 26.0 61.0 15.1 90.7 70.0 19.8 68.0 71.0 83.0 54.8 52.7
@@
Comparison on Representation Models We compare two 3D representation models in this experiment: TAP, which lacks rotation invariance, and TET, which is designed to address this issue. The comparison is conducted on the aligned-pose dataset ModelNet10 and the open-pose dataset ModelNet10. In addition, we only utilize Shap-e as the generative model. Experimental results, shown in Table 4, demonstrate that on the aligned-pose dataset, TAP and TET exhibit comparable performance. Specifically, TET’s oAcc and mAcc are merely 0.1% higher than those of TAP. This implies that the two models, despite being designed with distinct approaches, display minimal performance disparities on the aligned-pose dataset. Notably, on the open-pose dataset, TET significantly outperforms TAP across the majority of categories. For instance, TET shows a significant improvement over TAP’s oAcc of 21.5% and mAcc of 21.8%.
@@
It can be observed that designing specific structures to address point cloud rotation invariance effectively enhances the model’s robustness against rotation. The results show that TET, which incorporates such structures, is unaffected by point cloud rotations, as evidenced by the minimal difference in performance between the aligned-pose and open-pose datasets. In fact, TET performs slightly better on the open-pose dataset, with oAcc and mAcc increasing by 1.6% and 1.7%, respectively. In contrast, TAP, which uses a Transformer-based framework without specialized structures for rotation invariance, achieves results comparable to TET on the aligned-pose dataset. However, its performance drops sharply on the open-pose dataset, underscoring the need for rotation-invariant design to maintain robustness.
@@
Figure 4: Visualization comparison of ”dresser” point clouds from original category names ($\mathbf{CN}$) and ChatGPT-generated prompts ($\mathbf{CD}$), which enhance anchors’ diversity.
4.4 Discussion on LLM-Generated Category Descriptions
@@
During our experiments, we observed that the classification accuracy for certain categories remained consistently low. This was likely due to insufficient diversity in the generated samples. We hypothesized that generating more varied samples would improve the results. To address this, we explored optimizing descriptions for specific categories to improve classification performance. Taking the dresser category from the ModelNet10 dataset as an example, which had a notably low accuracy, we employed ChatGPT to generate multiple descriptive phrases tailored to the dresser category. These descriptions followed the format “a dresser that looks like … .” Using these optimized descriptions, we applied the Shap-e to create multiple new point clouds for the dresser category based on each description. The experimental results are presented in Tables 6, and the comparison between anchor points generated using ChatGPT-generated descriptions ($\mathbf{CD}$) versus category names ($\mathbf{CN}$) is shown in Fig. 4. The results demonstrate that appropriate descriptions significantly enhanced classification performance for the dresser category, while some inappropriate descriptions led to a decline in performance. However, since we did not know beforehand which descriptions would yield the best results, in subsequent experiments, we generated anchor samples using each description for all categories to further improve performance.
@@
We applied this approach on the ModelNet10 dataset, generating LLM-generated category descriptions to create anchor samples for classification tasks. From the results, shown in Table 7, the oAcc and mAcc were increased by 1.3% and 1.5%, respectively. However, for many individual categories, we observed that the performance was worse compared to the results generated using only class names. We believe this is due to the higher variance in the samples generated by the model when using ChatGPT-generated descriptions, which covered a broader range of features. However, this also caused interference between categories, such as generating a dresser that closely resembled a table. Due to the instability of the descriptions produced by LLMs, we did not include these results in comparisons with other methods. Hence, improving control over LLM-generated descriptions for better results is a promising future direction.
5 Conclusion
In this paper, we propose a pipeline utilizes 3D point cloud generative models for 3D open-world classification. Unlinke existing methods that rely on large-scale 2D priors, our approach directly leverages the prior knowledge from 3D generative models. This completely avoids the 3D-to-2D projection process, thus mitigating the instability in open-pose scenes that could degrade performance due to projection settings. We also explore how the number and diversity of generated anchor samples, as well as rotation-invariant feature spaces, affect the proposed pipeline’s performance.
Experiments demonstrate SOTA performance on open-world datasets, highlighting the proposed pipeline as a promising avenue for future research in 3D open-world classification.

### v2

@@ 3.4. Inference Phase
(5) $\hat{TS_{m}}=\frac{\mathbf{FPS}(TS_{m},\hat{N_{p}})-\mathbf{Avg}(TS_{m})}{\mathbf{L}(TS_{m})}\cdot R,\quad R\in\mathbb{R}^{3\times 3},$
@@
(6) $TF_{m}=\mathbf{RMs}\left(\hat{TS_{m}}\right),\quad TF_{m}\in\mathbb{R}^{\hat{N_{p}}\times D}.$
@@
4. Experiment
4.1. Experiment Setup
4.1.1. Datasets.
The effectiveness of the proposed method is validated on the public datasets ModelNet40, ModelNet10, and McGill, proposed in OP3D (Zhao et al., 2025), to assess performance under diverse geometric complexity. The ScanObjectNN (Uy et al., 2019) dataset, featuring real-world scanned data with occlusions and noise across 15 categories, is used to evaluate the effect of pretrained representation models.
@@ 4.1. Experiment Setup
ModelNet40/10 evolves from standard ModelNet40/10 (Wu et al., 2015) through random rotation of the samples, resulting in open-pose variations. ModelNet40 serves as a benchmark dataset within the domain of 3D point cloud classification. It originates from computer-aided design models and comprises high-quality 3D synthetic objects, which are devoid of noise and exhibit regular structural features. The dataset encompasses 40 categories of common objects, including pianos, tables, and airplanes, and contains a total of 12,311 3D objects, each represented by 10,000 points. These points are characterized by 3D coordinates (x, y, z) and may include additional features, such as normal vectors or color information (R, G, B). ModelNet10 is a specifically curated subset of ModelNet40, focusing on 10 common object categories, such as beds, monitors, and chairs, containing a total of 4,899 3D objects.
McGill (Zhao et al., 2025), an enhanced variant of the McGill dataset (Siddiqi et al., 2008), created through random sample rotation and the removal of five categories that overlap with ModelNet40. This modified dataset contains 14 distinct categories of complex geometric shapes (e.g., ants, snakes, and hands) characterized by significant morphological diversity. Natural object categories pose challenges for current methods due to complex topological variations.
@@
McGill (Zhao et al., 2025) is derived from McGill (Siddiqi et al., 2008) by randomly rotating the samples in McGill and removing five categories that overlap with ModelNet40. McGill includes 14 categories of complex geometric shapes, such as ants, snakes, and hands. The objects in this dataset exhibit higher morphological diversity, particularly posing greater challenges when handling natural object categories. Each sample consists of point coordinates (x, y, z) and topological information.
@@
Table 1. Comparison to SOTAs. Proj. indicates whether it is necessary to project the 3D samples into 2D images. denotes that there is an overlap between test categories and training categories of the pre-trained representation model utilized.
@@
Ours N 59.0 60.3 47.8 47.5
4.1.2. Evaluation Metric.
Following SOTAs (Zhang et al., 2023; Huang et al., 2023; Zhao et al., 2025), we evaluate the open-world classification performance using top-1 accuracy (oAcc) and mean accuracy (mAcc). oAcc measures the proportion of samples for which the model’s top prediction matches the ground truth, while mAcc calculates the average accuracy across all categories to assess the balance of performance among classes.
4.1.3. Implement Details.
All feature extraction and inference tasks are performed on an NVIDIA RTX 3090 GPU. For the point cloud diffusion model, we utilize a random seed $s\sim\mathcal{U}(0,50)$ to introduce diversity into the generated anchor point clouds of any category. During the data augmentation stage, all the anchor samples and test samples are downsampled to 1024 points by the farthest point sampling. For the feature output dimensions of different models, the TET pre-trained model outputs features with dimensions of (256, 1), and the TAP pre-trained model outputs features with dimensions of (256, 384). To facilitate subsequent feature comparison and processing, max pooling along TAP’s output second dimension unifies features to (256, 1).
Figure 3. Comparison on the Number of Anchor Samples $N_{a}$.
4.2. Performance Comparison
Our approach is compared against six SOTA methods: PointCLIP (Zhang et al., 2023), ULIP (Xue et al., 2023), ReconCLIP (Qi et al., 2023), CLIP2Point (Huang et al., 2023), PointCLIPv2 (Zhu et al., 2023), and OP3D (Zhao et al., 2025), on the open-pose classification datasets ModelNet10 and McGill. To ensure fairness, we utilize the pre-trained models from the official GitHub repositories provided by the respective papers (TAP (Wang et al., 2023) and TET (Melnyk et al., 2024)) and strictly followed the original experimental settings. For the tests on the McGill dataset, we employ a representation model trained on ModelNet40, as there is no category overlap between these datasets. However, due to the overlapping categories between ModelNet40 and ModelNet10, we manually exclude the overlapping categories from ModelNet40 to form ModelNet30, and subsequently trained a new representation model on this adjusted dataset for the ModelNet10 tests. The results presented in Section 4.2 and Section 4.3 do not include the samples generated from LLM-generated prompts. That is, all anchor samples are directly generated using the category names. The impact of LLM-generated prompts on the results will be discussed in subsequent sections.
Our approach exhibits notable enhancements in classification accuracy over current models in open-pose recognition tasks. The performance of our model on McGill dataset and ModelNet10 dataset are shown in Table 1. Notably, we achieved a 32.7% improvement in oAcc and a 36.1% improvement in mAcc on ModelNet10 dataset. On McGill dataset, our model also demonstrates competitive performance. Compared to the current leading method, our approach improve oAcc by 8.7% and mAcc by 2.8%.
Table 2. Comparison of classification performance using anchor samples generated by different generative models on randomly rotated dataset ModelNet10.
@@ 4.2. Performance Comparison
Shap-e 80.0 40.0 37.0 73.3 73.3 36.0 44.2 49.0 79.0 80.0 57.8 59.2
GD 74.0 34.0 47.0 75.6 76.7 38.0 23.3 58.0 66.0 79.0 56.2 57.2
Ensemble 80.0 27.0 43.0 77.9 74.4 39.0 38.4 67.0 77.0 79.0 59.0 60.3
Table 3. Comparison of classification performance using anchor samples generated by different generative models on randomly rotated dataset McGill. Abbreviations used: Din. = Dinosaur, Dol. = Dolphin, Oct. = Octopus, Qua. = Quadruple, Spe. = Spectacle, Spi. = Spider.
@@
Table 4. Comparative evaluation of classification performance under aligned-pose (ModelNet10) and open-pose (ModelNet10) settings across various representation models, with Shap-e used as the generative model. RMs denote the representation models used for extracting features.
@@
TET 74.0 47.0 47.0 74.4 68.6 44.0 54.7 34.0 73.0 85.0 59.1 60.2
(a) TET on aligned-pose
(b) TET on open-pose
(c) TAP on aligned-pose
(d) TAP on open-pose
Figure 4. t-SNE visualizations of feature representations extracted by TAP and TET under aligned-pose and open-pose settings on the ModelNet10 dataset. Different colors indicate different object categories according to the ground truth labels. Triangles represent generated anchor samples, while circles denote original dataset samples.
Table 5. Impact of TET pre-trained on different datasets when evaluated on ModelNet10
@@
ModelNet40 9840 40 10 90.2 90.3
ModelNet30 6049 30 0 59.0 60.3
ScanObjectNN 2309 15 7 54.6 52.0
4.3. Analysis of Results
The performance of the proposed method is influenced by several factors, including the number of anchor points, the selection of generative models, and 3D representation models. This section presents a quantitative analysis of each factor’s impact, showcasing our approach’s effectiveness and robustness.
4.3.1. Impact of Anchor Sample Size $N_{a}$
We examine the impact of varying the number of anchors generated while fixing the dataset, representation backbone (TET), and generation model to evaluate its effect on classification performance. Three distinct anchor generation methods are evaluated: Shap-e, GaussianDreamer (GD), and their mixed (Ensemble). As shown in Fig. 3, classification accuracy exhibits positive correlation with $N_{a}$, eventually reaching stable. However, on McGill, performance slightly decreased at 8 anchor samples, likely attributable to instability when anchor counts are low—where individual sample variance disproportionately influences feature statistics. This effect diminishes as $N_{a}$ increases. While more anchor samples generally improve performance, the marginal gain decreases beyond a point. This must be balanced against linearly increasing computational overhead during generation. Based on empirical convergence and computational trade-offs, we set $N_{a}=7$ as the optimal anchor points across all the experiments.
4.3.2. Comparison on Generative Models
The experiment maintains the constant dataset, representation model (TET), and number of anchor points, altering only the generative models to assess their impact on classification performance. Two generative models are selected: Shap-e and GD. Both models can generate point clouds based on the given prompts, but they differ in methodology. Shap-e directly generates a point cloud, whereas GD first generates a coarse point cloud and then optimizes it using 3D Gaussian Splatting. In terms of generation time, Shap-e requires 7 seconds per sample, while GD take 15 minutes per sample. The results on ModelNet10 dataset and McGill dataset are shown in Table 2 and Table 3, respectively. Overall, Shap-e outperforms GD in terms of classification accuracy on both datasets, but it does not consistently lead in every category. Both generative models have their strengths. On ModelNet10, Shap-e excels in some categories, e.g. bathtub, bed, and night stand, while GD performs better in others, e.g. chair, dresser, and sofa. On McGill, Shap-e performs significantly better in bird, hand, spectacle, and spider, while GD outperforms Shap-e in ant, dolphin, fish, and quadruped. We further conduct experiments by combining anchor samples generated by Shap-e and GD, referred to as Ensemble. The results demonstrate that the integration of anchor samples generated by both generative models results in enhanced overall performance. Significantly, the category-specific accuracy obtained by the Ensemble closely aligns with the leading generative model for each category. These results imply that the diversity inherent in samples produced by 3D generative models plays a beneficial role in augmenting the efficacy of the classification task.
Table 6. Comparison of prompts and the classification accuracies of samples generated for ‘dresser’.
Category Text Description $\mathbf{CD}$ oAcc of ‘Dresser’$\uparrow$
A dresser. 73.3
A dresser with five drawers and circular handles. 68.6
An open design dresser with five exposed drawers. 72.1
A small bedside dresser with two drawers”. 68.6
A vintage dresser with carved details and two doors. 76.7
A decorative dresser with floral carvings and four drawers. 57.0
@@ 4.3.2. Comparison on Generative Models
Table 7. Comparison of classification result on ModelNet10 using Shap-e with Category Name ($\mathbf{CN}$) and ChatGPT ($\mathbf{CD}$) prompts.
@@
$\mathbf{CN}$ 80.0 40.0 37.0 73.3 73.3 36.0 44.2 49.0 79.0 80.0 57.8 59.2
$\mathbf{CD}$ 82.0 44.0 50.0 74.4 72.1 31.0 43.0 43.0 82.0 86.0 59.5 60.8
4.3.3. Comparison on Representation Models
@@
We compare two 3D representation models in this experiment: TAP, which lacks rotation invariance, and TET, which is designed to address this issue. The comparison is conducted on the aligned-pose dataset ModelNet10 and the open-pose dataset ModelNet10. In addition, we only utilize Shap-e as the generative model. Experimental results, shown in Table 4, demonstrate that on the aligned-pose dataset, TAP and TET exhibit comparable performance. This implies that the two models, despite being designed with distinct approaches, display minimal performance disparities on the aligned-pose dataset. Notably, on the open-pose dataset, TET significantly outperforms TAP across the majority of categories. For instance, TET shows a significant improvement over TAP’s oAcc of 36.3% and mAcc of 37.4%.
@@
It can be observed that designing specific structures to address point cloud rotation invariance effectively enhances the model’s robustness against rotation. The results show that TET, which incorporates such structures, is unaffected by point cloud rotations, as evidenced by the minimal difference in performance between the aligned-pose and open-pose datasets. In fact, TET exhibits a slight performance drop on the open-pose dataset, with oAcc and mAcc decreasing by 1.3% and 1.0%, respectively. In contrast, TAP, which uses a Transformer-based framework without specialized structures for rotation invariance, achieves results comparable to those of TET on the aligned-pose dataset. However, its performance drops sharply on the open-pose dataset, underscoring the need for rotation-invariant design to maintain robustness.
To provide a more intuitive understanding of this issue, as shown in Fig. 4, we visualize the feature representations of TAP and TET on the ModelNet10 dataset using t-SNE under both aligned-pose and open-pose settings. In the aligned-pose scenario, TAP and TET exhibit similarly well-clustered feature distributions. The generated samples (marked with triangles) also contribute effectively to category separation. However, under the open-pose setting, TAP’s t-SNE plot becomes cluttered and disorganized, indicating that its features are less distinguishable due to sensitivity to rotation. In contrast, TET maintains clear and compact clustering, further confirming its robustness to arbitrary rotations.
4.3.4. Effect of Representation Models on Pre-trained Datasets
We study the impact of pre-training representation models for open-world classification by assessing the TET model, pre-trained on ModelNet40, ModelNet30, and ScanObjectNN, on the ModelNet10 dataset, as illustrated in Table 5. When the pre-training dataset fully overlaps with the test categories (i.e., ModelNet40 includes all ModelNet10 categories), the model achieves an oAcc of 90.2%, a result competitive with those obtained via fully supervised learning. This demonstrates that our generated samples contain valid and discriminative category-level information, and also confirms the effectiveness of the proposed pipeline. However, when forming ModelNet30 by removing overlapping categories, performance drops significantly, indicating the challenge category-level openness poses, even for rotation-invariant models. Furthermore, using ScanObjectNN, with fewer categories and samples and a distinct data distribution (real-world scans instead of CAD models), the model still achieves an oAcc of 54.6% despite 7 overlapping categories, indicating that out-of-distribution (OOD) issues persist in rotation-invariant models. On the other hand, the result highlights that high category overlap alone does not guarantee high performance when the pre-training and testing distributions differ significantly.
@@ 4.3.4. Effect of Representation Models on Pre-trained Datasets
Figure 5. Visualization comparison of ”dresser” point clouds from original category names ($\mathbf{CN}$) and ChatGPT-generated prompts ($\mathbf{CD}$), which enhance anchors’ diversity.
4.4. Discussion on LLM-Generated Category Descriptions
@@
In our experiments, some categories have consistently low classification accuracy, likely due to a lack of sample diversity. We hypothesize that more varied samples could enhance results by optimizing descriptions for specific categories. Taking the dresser category from the ModelNet10 dataset as an example, which has notably low accuracy, we employ ChatGPT to generate multiple descriptive phrases tailored to the dresser category. These descriptions followed the format “a dresser that looks like … .” Using these optimized descriptions, we apply the Shap-e to create multiple new point clouds for the dresser category based on each description. The experimental results are presented in Tables 6, and the comparison between anchor points generated using ChatGPT-generated descriptions ($\mathbf{CD}$) versus category names ($\mathbf{CN}$) is shown in Fig. 7. The results demonstrate that appropriate descriptions significantly enhanced classification performance for the dresser category, while some inappropriate descriptions led to a decline in performance. However, since we did not know beforehand which descriptions would yield the best results, in subsequent experiments, we generated anchor samples using each description for all categories to further improve performance.
@@
We apply this approach on the ModelNet10 dataset, generating LLM-generated category descriptions to create anchor samples for classification tasks. From the results, shown in Table 7, the oAcc and mAcc are increased by 1.7% and 1.6%, respectively. However, for many individual categories, we observe that the performance is worse compared to the results generated using only class names. The observed performance fluctuations are more likely due to differences in how well the prompts align with the core visual features of the target category. However, this also causes interference between categories, such as generating a dresser that closely resembles a table. We exclude these results from method comparisons due to the instability of LLM-generated descriptions. Hence, improving control over LLM-generated descriptions for better results is a promising future direction.
5. Conclusion
This paper proposes a pipeline that leverages 3D point cloud generative models for open-world 3D classification. Unlike existing methods that rely on 2D priors and projection, our approach directly utilizes prior knowledge from 3D generative models, avoiding projection-induced instability in open-pose scenarios. The method is training-free, rotation-invariant, and capable of adapting to novel categories with only a few generated samples—an advantage especially valuable given the high cost of collecting real-world 3D data.
We further examine how anchor sample quantity, diversity, and feature space design affect performance, and explore the role of LLM-generated prompts. Performance variability across prompts is found to stem not from linguistic ambiguity, but from differences in how well prompts capture the visual and structural essence of the target category. Future work may explore improving prompt strategies to generate more diverse and representative samples that better reflect core category features.